an unfair coin lands heads 65% of the time. This coin is flipped 1500 times. What is the probability that the coin will land heads between 960 and 1000 times
Answer by Boreal(15235)

the expected value is 1500*0.65=975 times=np
the variance is np(1-p)=975*0.35=341.25
the sd is the sqrt (341.25)=18.47
So, this can be approximated by a normal distribution
z=(960-975)/18.47=-0.812
z=(1000-975)/18.47=1.353
The probability is that part between these two z-values
it is 70.36% or 0.7036